OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (radar):

Part 1: First do you know what the form the "point-slope form" is? here is the form: y1-y=m(x1-X) where m = slope. You have been given x1 and y1 as well as the slope. Plug in the values and write the equation. I will standby for questions and problems that you may incur.

OpenStudy (radar):

-6-y=4(3-x) -6-y=12-4x -y=-4x+18 y=4x-18 slope intercept form Part 2

OpenStudy (radar):

y-4x=-18 Part 3 (subtracted (4x) from both sides.